OpenStudy (anonymous):

Solve -[(2y+4)-(3y-7)]-(5-y)

OpenStudy (snowfire):

Combine like terms and simplify. Be careful with distributing the negative.

hero (hero):

\(-[(2y+4)-(3y-7)]-(5-y)=-[2y+4-3y+7]-(5-y)\) \(=-[2y-3y+4+7]-(5-y)\) \(=-[-y+11]-(5-y)\) \(=y-11+y-5\) \(=y+y-11-5\) \(=2y-16\)